# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ III.2 Frequency-domain PIRD
To solve the PIRD identified by (13), it is necessary to define a redundancy rate function taking values over the lattice underlying the decomposition. In principle, the redundancy rate can be defined following any of the several approaches formulated for the PID, adapting it to the calculation of the MIR between random processes in place of the MI between random variables. In the PID literature, after the seminal work by Williams and Beer who first recognized that the concept of redundancy cannot be defined using classical information theory and proposed an axiomatic definition of redundant information Williams and Beer (2010), a number of alternative redundancy measures adopting or modifying the original set of axioms have been proposed Bertschinger et al. (2014); Barrett (2015); Ince (2017); Pakman et al. (2021); Makkeh et al. (2021); Ehrlich et al. (2024). A particularly useful approach is to derive redundancy measures working on the specific realizations of the random variables at hand, rather than on the variable themselves, and then compute redundancy via statistical expectation. This pointwise approach has been followed to put forth PID methods that define "local" redundancy measures computed for single realizations of the variables, from which a "global" redundancy is obtained taking the ensemble average over all possible realizations Ince (2017); Finn and Lizier (2018); Makkeh et al. (2021); Gutknecht et al. (2021). The same approach can be followed for the PIRD, e.g., working on the local version of the MIR (12) to formalize the notion of pointwise redundancy rate; the redundancy rate among processes could be then retrieved by ensemble averaging, which for stationary processes corresponds to time-domain averaging.
@@
Here, we propose an approach that is conceptually similar, but is implemented through a pointwise representation in frequency rather than in time. Specifically, we characterize the analyzed network of random processes in the frequency domain, considering the information provided about a particular oscillatory component of the target process $Y$ by the iso-frequency oscillatory components of the source processes collected in $X$. The idea is to perform the entire PIRD on the pointwise level for a particular frequency, i.e., to decompose the spectral (frequency-specific) MIR denoted as $i_{Y;X}(\omega)$, where $\omega=2\pi\frac{f}{f_{s}}$ is the normalized circular frequency ($\omega\in[-\pi,\pi]$, $f\in[-\frac{f_{s}}{2},\frac{f_{s}}{2}]$); the spectral MIR is identified from the expansion of the MIR in the frequency domain as:
$I_{Y;X}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i_{Y;X}(\omega)\mathrm{d}\omega.$ (17)
@@
While different integral transforms (e.g., the wavelet transform) could in principle be used to expand the MIR, in the next section we will use a definition of spectral MIR which satisfies (17) for Gaussian processes to formalize the PIRD in the frequency domain. Crucially, (17) connects the time- and frequency-domain representations of information-theoretic quantities for random processes, and is exploited here to relate the PIRD (13) to its frequency-domain extension. Such an extension is denoted as spectral PIRD and is formulated, for the oscillatory components of the network process assessed at the frequency $\omega$, expressing the spectral MIR as:
$i_{Y;X}(\omega)=\sum_{\alpha\in\mathcal{A}}i^{\delta}_{Y;X_{\alpha}}(\omega),$ (18)
@@
where $i^{\delta}_{\cdot;\cdot}(\omega)$ is the spectral partial information rate function defined over a lattice specifically identified on the oscillations with frequency $\omega$, and $X_{\alpha}$ denotes the set of subsets of source processes indexed by the atom $\alpha$. As happens with PID, to solve (18) we need to identify a spectral redundancy rate through the imposition of a certain number of reasonable axioms or constraints on the spectral MIR, and then follow the PID formalism to derive the pointwise (spectral) atoms of information at the frequency $\omega$. If $i^{\cap}_{Y;X_{\alpha}}(\omega)$ denotes the spectral redundancy rate computed for the $\alpha^{\mathrm{th}}$ node of the spectral lattice defined at the frequency $\omega$, the corresponding spectral PI rate is computed recursively, in analogy to (5), as:
$i^{\delta}_{Y;X_{\alpha}}(\omega)=i^{\cap}_{Y;X_{\alpha}}(\omega)-\sum_{\beta\prec\alpha}i^{\delta}_{Y;X_{\beta}}(\omega).$ (19)
@@
$i_{Y;X}(\omega)=\sum_{m=1}^{M}u_{Y;X_{m}}(\omega)+r_{Y;X}(\omega)+s_{Y;X}(\omega),$ (20)
@@
Eq. (19) provides a solution for the spectral PIRD up to the definition of a proper spectral redundancy rate function. Here, we propose to assess spectral redundancy following the minimum mutual information (MMI) principle Barrett (2015) applied to the spectral MIR computed between the target and each source process at the frequency of interest. Specifically, we define the frequency-specific redundancy rate function of the atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$ of the spectral redundancy lattice as:
$i^{\cap}_{Y;X_{\alpha}}(\omega)=\min\limits_{j=1,\ldots,J}i_{Y;X_{\alpha_{j}}}(\omega),$ (21)
@@
where $J=|\alpha|$ indicates the cardinality of the atom; e.g., $J=1$ if $\alpha=\{1\}$ or $\alpha=\{12\}$, while $J=2$ if $\alpha=\{\{3\},\{12\}\}$. Note that, in contrast with the classical MMI formulation performed for random variables Barrett (2015), here the minimum MIR is searched at the pointwise (frequency-specific) level. While a pointwise implementation of the MMI criterion would be cumbersome if performed in the time domain, and it is indeed avoided by the existing pointwise approaches Ince (2017); Finn and Lizier (2018); Makkeh et al. (2021) because the local MI or the local MIR can take negative values, in our case is favored by the non-negativity of the spectral MIR. The properties of the spectral redundancy rate (21) computed for Gaussian processes will be discussed in the following subsection.
@@
Importantly, from the perspective of information decomposition based on lattice structures Williams and Beer (2010), the spectral PIRD (18,19) is conceptually equivalent to the PIRD (13,14) and to the PID (3,4); what changes in the three formulations is only the quantity to be decomposed, from the MI $I(T;S)$ for the case of random variables to the MIR $I_{Y;X}$ for the case of random processes, and to the spectral MIR $i_{Y;X}(\omega)$ for the case of oscillatory components of random processes. Since these quantities maintain the same meaning and properties (i.e., they are non-negative measures of shared information), the three formulations of PID, PIRD and spectral PIRD will lead to a unique solution for the atoms once a redundancy function (respectively, the redundancy among collections of random variables, the redundancy rate among collections of random processes, and the spectral redundancy rate among collections of oscillatory components of random processes) is fixed over a lattice structure like that in Fig. 1. Therefore, moving from the PIRD defined on the pointwise level (frequency-specific) to that defined on the process-level (time domain) is straightforward exploiting spectral integration applied to the redundancy rate and PI rate functions:
$I^{\cap}_{Y;X_{\alpha}}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i^{\cap}_{Y;X_{\alpha}}(\omega)\mathrm{d}\omega,$ (22)
$I^{\delta}_{Y;X_{\alpha}}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i^{\delta}_{Y;X_{\alpha}}(\omega)\mathrm{d}\omega;$ (23)
the same spectral integration property holds to relate the unique, redundant and synergistic atoms of the coarse-grained spectral PIRD (20) to their corresponding time-domain atoms of the PIRD (15).
Remarkably, solving the frequency-specific PIRD to obtain the PI rate $i^{\delta}_{Y;X_{\alpha}}(\omega)$ for each atom $\alpha$ via (19) and then integrating these contributions along the whole frequency axis according to (23) to get time domain values, is equivalent to integrating the spectral redundancy rate functions $i^{\cap}_{Y;X_{\alpha}}(\omega)$ in the range $[-\pi,\pi]$ via (22) and then applying the PIRD in the time domain to get the contributions $I^{\delta}_{Y;X_{\alpha}}$. This is guaranteed by the property of linearity of the definite integrals, and allows also to develop band-specific PIRD schemes whereby the decomposition of multivariate information rates is achieved in the time domain but is limited to oscillatory components whose frequencies are confined within a specific band of the spectrum. This band-limited decomposition, which is useful to highlight and decompose multivariate interactions with focus on specific oscillations with practical meaning, will be illustrated in the next sections in both simulated and applicative settings.
@@ III.3 Formulation for Gaussian processes
In the linear signal processing framework, the analyzed set of stochastic processes $Z=\{Y,X_{1},\ldots,X_{M}\}$ can be described in terms of its power spectral density (PSD) matrix expressed as
$\textbf{P}_{Z}(\omega)=\begin{bmatrix}P_{Y}(\omega)&P_{YX_{1}}(\omega)&\cdots&P_{YX_{M}}(\omega)\\ P_{X_{1}Y}(\omega)&P_{X_{1}}(\omega)&\cdots&P_{X_{1}X_{M}}(\omega)\\ \vdots&\vdots&\ddots&\vdots\\ P_{X_{M}Y}(\omega)&P_{X_{M}X_{1}}(\omega)&\cdots&P_{X_{M}}(\omega)\\ \end{bmatrix},$ (24)
which contains the individual PSDs of the processes $\{Y,X_{1},\ldots,X_{M}\}$ as diagonal elements and the cross-PSDs between them as off-diagonal elements. Individual and cross-PSDs are defined as the Fourier Transform (FT) of the covariance functions of the processes: i.e., $P_{Y}(\omega)=\mathfrak{F}\{R_{Y}(k)\}$, $P_{X_{i}X_{j}}(\omega)=\mathfrak{F}\{R_{X_{i}X_{j}}(k)\}$ and $P_{YX_{i}}(\omega)=\mathfrak{F}\{R_{YX_{i}}(k)\}$, where $R_{Y}(k)=\mathbb{E}[Y_{n}Y_{n-k}]$, $R_{X_{i}X_{j}}(k)=\mathbb{E}[X_{i,n}X_{j,n-k}]$ and $R_{YX_{i}}(k)=\mathbb{E}[Y_{n}X_{i,n-k}]$ are the auto-correlation function of $Y$, the cross-correlation function between $X_{i}$ and $X_{j}$ and the cross-correlation function between $Y$ and $X_{i}$ ($k$ is the correlation delay), respectively; correlations are equivalent to covariances for zero-mean processes.
@@
$I_{X;Y}=\frac{1}{4\pi}\int_{-\pi}^{\pi}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log}\frac{|\textbf{P}_{X}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{Z}(\omega)|}\mathrm{d}\omega,$ (25)
where $\textbf{P}_{X}(\omega)$ is the PSD matrix pruned from the first row and column, and ||$\cdot$|| ⋅ | denotes matrix determinant. The spectral MIR decomposed by PIRD then results immediately comparing (25) with (17):
$i_{Y;X}(\omega)=\frac{1}{2}\log\frac{|\textbf{P}_{X}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{Z}(\omega)|}.$ (26)
@@
$i_{Y;X_{\alpha_{j}}}(\omega)=\frac{1}{2}\log\frac{|\textbf{P}_{X_{\alpha_{j}}}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{[YX_{\alpha_{j}}]}(\omega)|},$ (27)
where
$\textbf{P}_{[YX_{\alpha_{j}}]}(\omega)=\begin{bmatrix}P_{Y}(\omega)&\textbf{P}_{YX_{\alpha_{j}}}(\omega)\\ \textbf{P}_{X_{\alpha_{j}}Y}(\omega)&\textbf{P}_{X_{\alpha_{j}}}(\omega)\\ \end{bmatrix};$ (28)

### v3

@@ III.2 Frequency-domain PIRD
To solve the PIRD identified by (13), it is necessary to define a redundancy rate function taking values over the lattice underlying the decomposition. In principle, the redundancy rate can be defined following any of the several approaches formulated for the PID, adapting it to the calculation of the MIR between random processes in place of the MI between random variables. In the PID literature, after the seminal work by Williams and Beer who first recognized that the concept of redundancy cannot be defined using classical information theory and proposed an axiomatic definition of redundant information Williams and Beer (2010), a number of alternative redundancy measures adopting or modifying the original set of axioms have been proposed Bertschinger et al. (2014); Barrett (2015); Ince (2017); Pakman et al. (2021); Makkeh et al. (2021); Ehrlich et al. (2024). Among them, a simple definition of redundancy is the MMI measure Barrett (2015), which can be straightforwardly adapted to information rates; specifically, a so-called MMI redundancy rate can be defined taking the minimum MIR over the source processes composing the analyzed atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$ of the redundancy lattice:
$I^{\cap}_{Y;\textbf{X}_{\alpha}}=I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}:=\min\limits_{j=1,\ldots,J}I_{Y;\textbf{X}_{\alpha_{j}}}.$ (17)
@@
As an alternative to using directly (17), in this work we exploit the MMI principle at the pointwise level in the spectral domain. Traditional pointwise methods applied to multiple random variables derive "local" redundancy measures working on the specific realizations of the variables at hand, rather than on the variable themselves, and then compute "global" redundancy via statistical expectation, i.e. taking the ensemble average over all possible realizations Ince (2017); Finn and Lizier (2018); Makkeh et al. (2021); Gutknecht et al. (2021). The same approach can be followed for the PIRD, e.g., working on the local version of the MIR (12) to formalize the notion of pointwise redundancy rate; the redundancy rate among processes could be then retrieved by ensemble averaging, which for stationary processes corresponds to time-domain averaging. Here, we propose an approach that is conceptually similar, but is implemented through a pointwise representation in frequency rather than in time. Specifically, we characterize the analyzed network of random processes in the frequency domain, considering the information provided about a particular oscillatory component of the target process $Y$ by the iso-frequency oscillatory components of the source processes collected in X. The idea is to perform the entire PIRD on the pointwise level for a particular frequency, i.e., to decompose the spectral (frequency-specific) MIR denoted as $i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)$, where $\omega=2\pi\frac{f}{f_{s}}$ is the normalized circular frequency ($\omega\in[-\pi,\pi]$, $f\in[-\frac{f_{s}}{2},\frac{f_{s}}{2}]$); the spectral MIR is identified from the expansion of the MIR in the frequency domain as:
$I_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\frac{1}{2\pi}\int_{-\pi}^{\pi}i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)\mathrm{d}\omega.$ (18)
@@
While different integral transforms (e.g., the wavelet transform) could in principle be used to expand the MIR, in the next section we will use a definition of spectral MIR which satisfies (18) for Gaussian processes to formalize the PIRD in the frequency domain. Crucially, (18) connects the time- and frequency-domain representations of information-theoretic quantities for random processes, and is exploited here to relate the time-domain PIRD (13) to its frequency-domain extension. Such an extension is denoted as spectral PIRD and is formulated, for the oscillatory components of the network process assessed at the frequency $\omega$, expressing the spectral MIR as:
$i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)=\sum_{\alpha\in\mathcal{A}}i^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega),$ (19)
@@
where $i^{\delta}_{\cdot;\cdot}(\omega)$ is the spectral partial information rate function defined over a lattice specifically identified on the oscillations with frequency $\omega$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}$ denotes the set of subsets of source processes indexed by the atom $\alpha$. As happens with PID, to solve (19) we need to identify a spectral redundancy rate, $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)$, generalizing the spectral MIR over the lattice, and express it as the sum of the spectral PI rates of the atoms positioned downwards in the lattice:
$i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)=\sum_{\beta\preceq\alpha}i^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\beta}}(\omega).$ (20)
From (20), the spectral PI rate can be computed recursively, in analogy to (5), as:
$i^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)=i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)-\sum_{\beta\prec\alpha}i^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\beta}}(\omega).$ (21)
@@
$i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)=\sum_{m=1}^{M}u_{Y;X_{m}}(\omega)+r_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)+s_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega),$ (22)
@@
Eq. (21) provides a solution for the spectral PIRD up to the definition of a proper spectral redundancy rate function. Here, we propose to assess spectral redundancy following the MMI principle Barrett (2015) applied to the spectral MIR computed between the target and each source process at the frequency of interest. Specifically, we define the frequency-specific redundancy rate function of the atom $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$ of the spectral redundancy lattice, denoted as spectral minimum mutual information (SMMI) redundancy rate, as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i^{\cap}_{Y;\textbf{X}_{\alpha}}(\omega)=i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega):=\min\limits_{j=1,\ldots,J}i_{Y;\textbf{X}_{\alpha_{j}}}(\omega),}$ (23)
@@
where $J=|\alpha|$ indicates the cardinality of the atom; e.g., $J=1$ if $\alpha=\{1\}$ or $\alpha=\{12\}$, while $J=2$ if $\alpha=\{\{3\},\{12\}\}$. Note that, in contrast with the classical MMI formulation performed for random variables Barrett (2015), here the minimum MIR is searched at the pointwise (frequency-specific) level. While a pointwise implementation of the MMI criterion would be cumbersome if performed in the time domain, and it is indeed avoided by the existing pointwise approaches Ince (2017); Finn and Lizier (2018); Makkeh et al. (2021) because the local MI or the local MIR can take negative values, in our case is favored by the non-negativity of the spectral MIR. The properties of the spectral redundancy rate (23) computed for Gaussian processes will be discussed in the following subsection and in Appendix A.
Eq. (23) defines the redundancy rate and the related PIRD in the frequency domain. Then, to quantify the redundancy rate in the time domain we exploit the idea pursued by pointwise approaches of averaging local redundancy measures to get global measures Ince (2017); Finn and Lizier (2018); Makkeh et al. (2021); Gutknecht et al. (2021), with the difference that averaging is performed over frequency rather than over time. Specifically, the time-domain (global) version of the SMMI redundancy rate is defined as the normalized full-frequency integral of the spectral redundancy:
$I^{\cap}_{Y;\textbf{X}_{\alpha}}=I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}:=\frac{1}{2\pi}\int_{-\pi}^{\pi}i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega)\mathrm{d}\omega.$ (24)
The redundancy rate defined in (24) is finally used to derive the PI rate associated to the $\alpha^{\mathrm{th}}$ atom of the lattice in analogy to (5) as:
$I^{\delta}_{Y;\textbf{X}_{\alpha}}=I^{\cap}_{Y;\textbf{X}_{\alpha}}-\sum_{\beta\prec\alpha}I^{\delta}_{Y;\textbf{X}_{\beta}},$ (25)
or in a compact way via Möbius inversion of (14), thus completing the PIRD. Moreover, the PI rates of the various atoms can be aggregated as in (15) to yield a coarse-grained PIRD. Importantly, the SMMI redundancy rate defined in (24) is different than the MMI redundancy rate given in (17); in Appendix A.1 we demonstrate that $I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}\leq I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}$, with equivalence holding in the absence of temporal correlations (i.e. when the processes reduce to i.i.d. random variables). This discrepancy occurs because the MMI definition is not additive in the sense that averaging over frequencies the minimum spectral MIR does not yield the minimum time-domain MIR. Nevertheless, we take this as an advantage of the definition of redundancy rate (24), as a less conservative measure of redundancy can help reducing the overestimation of redundancy in which definitions based on the minimum MI principle may incur Timme et al. (2014); Griffith and Ho (2015); Harder et al. (2013); Ince (2017).
@@
From the perspective of information decomposition based on lattice structures Williams and Beer (2010), the PID (2,4) is conceptually equivalent to the PIRD (13,14) and to the spectral PIRD (19,20); what changes in the three formulations is only the quantity to be decomposed, from the MI $I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})$ for the case of random variables to the MIR $I_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ for the case of random processes, and to the spectral MIR $i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)$ for the case of oscillatory components of random processes. Since these quantities maintain the same meaning and properties (i.e., they are non-negative measures of shared information), the three formulations of PID, PIRD and spectral PIRD will lead to a unique solution for the atoms once a redundancy function (respectively, the redundancy among collections of random variables, the redundancy rate among collections of random processes, and the spectral redundancy rate among collections of oscillatory components of random processes) is fixed over a lattice structure like that in Fig. 1. Moreover, moving from the PIRD defined on the pointwise level (frequency-specific) to that defined on the process-level (time domain) is straightforward exploiting spectral integration. This is guaranteed by the property of linearity of the definite integrals, which allows e.g. to obtain (25) through full-frequency integration of (21). In fact, solving the frequency-specific PIRD to obtain the PI rate $i^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)$ for each atom $\alpha$ via (21) and then integrating these contributions along the whole frequency axis to get time domain values is equivalent to integrating the spectral redundancy rate functions $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)$ in the range $[-\pi,\pi]$ via (24) and then applying (25) in the time domain to get the contributions $I^{\delta}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}$ (proof in the Appendix A.2). The tight link between the spectral and time-domain formulations of the proposed PIRD schemes allows also the useful possibility of tailoring the information decomposition to predetermined frequency bands with practical meaning. Indeed, frequency-specific PIRD schemes can be designed whereby the decomposition of multivariate information rates is achieved in the time domain but is limited to oscillatory components whose frequencies are confined within a specific band of the spectrum. This band-limited decomposition, which is useful to highlight and decompose multivariate interactions with focus on specific oscillatory components of the analyzed processes, will be illustrated in the next sections in both simulated and applicative settings.
@@ III.3 Formulation for Gaussian processes
In the linear signal processing framework, the analyzed set of stochastic processes ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}=\{Y,X_{1},\ldots,X_{M}\}$ can be described in terms of its power spectral density (PSD) matrix expressed as
$\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}}(\omega)=\begin{bmatrix}P_{Y}(\omega)&P_{YX_{1}}(\omega)&\cdots&P_{YX_{M}}(\omega)\\ P_{X_{1}Y}(\omega)&P_{X_{1}}(\omega)&\cdots&P_{X_{1}X_{M}}(\omega)\\ \vdots&\vdots&\ddots&\vdots\\ P_{X_{M}Y}(\omega)&P_{X_{M}X_{1}}(\omega)&\cdots&P_{X_{M}}(\omega)\\ \end{bmatrix},$ (26)
which contains the individual PSDs of the processes $\{Y,X_{1},\ldots,X_{M}\}$ as diagonal elements and the cross-PSDs between them as off-diagonal elements. Individual and cross-PSDs are defined as the Fourier Transform (FT) of the covariance functions of the processes: i.e., $P_{Y}(\omega)=\mathfrak{F}\{R_{Y}(k)\}$, $P_{X_{i}X_{j}}(\omega)=\mathfrak{F}\{R_{X_{i}X_{j}}(k)\}$ and $P_{YX_{i}}(\omega)=\mathfrak{F}\{R_{YX_{i}}(k)\}$, where $R_{Y}(k)=\mathbb{E}[Y(t_{n})Y(t_{n-k})]$, $R_{X_{i}X_{j}}(k)=\mathbb{E}[X_{i}(t_{n})X_{j}(t_{n-k})]$ and $R_{YX_{i}}(k)=\mathbb{E}[Y(t_{n})X_{i}(t_{n-k})]$ are the auto-correlation function of $Y$, the cross-correlation function between $X_{i}$ and $X_{j}$ and the cross-correlation function between $Y$ and $X_{i}$, respectively ($k$ is the correlation delay); correlations are equivalent to covariances for zero-mean processes.
@@
$I_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}};Y}=\frac{1}{4\pi}\int_{-\pi}^{\pi}\log\frac{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}}(\omega)|}\mathrm{d}\omega,$ (27)
where $\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)$ is the PSD matrix pruned of the first row and column, and ||$\cdot$|| ⋅ | denotes matrix determinant. The spectral MIR decomposed by the spectral PIRD then results immediately comparing (27) with (18):
$i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)=\frac{1}{2}\log\frac{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}}(\omega)|}.$ (28)
@@
$i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)=\frac{1}{2}\log\frac{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{[Y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}]}(\omega)|},$ (29)
where $\textbf{P}_{[Y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}]}(\omega)$ is a sub-matrix identified from $\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}}(\omega)$ as
$\textbf{P}_{[Y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}]}(\omega)=\begin{bmatrix}P_{Y}(\omega)&\textbf{P}_{Y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)\\ \textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}Y}(\omega)&\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)\\ \end{bmatrix};$ (30)